OpenStudy (anonymous):

Evaluate the following expression using the values given: @tootles please help Find -a2 - 3b3 + c2 + 2b3 - c2 if a = 3, b = 2, and c = -3.

OpenStudy (anonymous):

@Tootles143 please help

OpenStudy (anonymous):

thankyou orange do u know this stuff

OpenStudy (anonymous):

all you have to do is replace the letters with their corresponding numbers so -3*2 - 3*2*3 + -3*2 (I'll leave the rest to you ) 2b3 - c2

OpenStudy (anonymous):

Hope I helped

OpenStudy (anonymous):

so is the answer -31

OpenStudy (anonymous):

1 sec I'll check

OpenStudy (anonymous):

thnk u

OpenStudy (anonymous):

not quite

OpenStudy (anonymous):

I have answer choices would u like to see them

OpenStudy (anonymous):

keep careful track of the - signs remember that - a negative is the same as + a posotive

OpenStudy (anonymous):

Evaluate the following expression using the values given: Find -a2 - 3b3 + c2 + 2b3 - c2 if a = 3, b = 2, and c = -3. -67 -17 -31 1

OpenStudy (anonymous):

ok i'll check my answer one more time looks like i did something wrong haha

OpenStudy (anonymous):

thankyou it looks more like this though

OpenStudy (anonymous):

ooooooohhhhh

OpenStudy (anonymous):

those are exponents not multiples

OpenStudy (anonymous):

yeah its confusing

OpenStudy (anonymous):

you should write it "-a^2 - 3b^3 + c^2 + 2b^3 - c^2"

OpenStudy (anonymous):

so to work it out all you have to do is know how exponents work

OpenStudy (anonymous):

yeah so would it be -31

OpenStudy (anonymous):

i think

OpenStudy (anonymous):

hbu

OpenStudy (anonymous):

i'll check haha was checking the wrong problem

OpenStudy (anonymous):

ok so i'll tell you that you're wrong but let's work out why

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

-3^2=-6 right

OpenStudy (anonymous):

so first we can say that we have -a^2 - 3b^3 + 2b^3 because c^2 and - c^2 cancel out alright?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so then we replace the letters with the numbers they stand for - 3^2 - 3*2^3 + 2*2^3

OpenStudy (anonymous):

now do you know how "^" work?

OpenStudy (anonymous):

basically if it says ^2 it means that there are two of that number being multiplied together

OpenStudy (anonymous):

yes thank u so it would be -67

OpenStudy (anonymous):

so 2^2 = 2*2 = 4 3^2 = 3*3 = 9 and 2^3 = 2*2*2 = 8

OpenStudy (anonymous):

fill it in like that and come back with your answer :)

OpenStudy (anonymous):

remember you're working from - 3^2 - 3*2^3 + 2*2^3

OpenStudy (anonymous):

u still there

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

so all you have to do is this -3*3 - 3*2*2*2 + 2*2*2*2

OpenStudy (anonymous):

-17

OpenStudy (anonymous):

Congratz

OpenStudy (anonymous):

thnks can u help with another please

OpenStudy (anonymous):

ok but i'm starting my own work so i might be a bit short with my responces

OpenStudy (anonymous):

ok

OpenStudy (deana99):

message me please